Find the molarity of individual ions in the following solution : 44.0mL of 0.35M NaCl and 25.0 mL of 0.24M CaCl2 are mixed together.
Answer #1

Given Nach 44ml, 0.35M Cacl5 25ml, 0.2419 we know that number of moles = concentration x Volume (int) n = Miu (inn) Nacht NacNaci Nat +cí cach - cat? + aci In the solution total ions are 1 nat, 7 cath, a ci Nach Nat + ci 1-54800 a mol so wat = ci = 1| c ] moles! 1.54x 10 mol + 12x10-3 mol (1.54 + 1-2)x162 mol = 2.74x16 mol 2.74x10-7 mol 69x103 - :: M = - 0.397 M .(CT) = 0.